# Deep SPI: Safe Policy Improvement via World Models

### v1

@@ 7 Experiments
In this section, we evaluate the practical performance of DeepSPI in environments where (i) representation learning is essential and (ii) dynamics are complex. We use the Atari Arcade Learning Environment (ALE; Bellemare et al. 2013) and represent each state by four stacked frames. Although ALE domains feature diverse dynamics, they are largely deterministic. To introduce stochasticity, we follow Machado et al. (2018) and employ two standard tricks: sticky actions, where with probability $p_{a}$ the previous action is repeated (simulating joystick or reaction-time imperfections), and random initialisation, where the agent begins after $n_{\text{NOOP}}$ initial no-op frames. We set $p_{a}=0.3$ and $n_{\text{NOOP}}=60$.
Figure 3: Relative improvement of DeepSPI compared to PPO over the full stochastic ALE suite (41/61).
Figure 4: Human normalized score over the stochastic, standard 57 envs. from ALE. Plots per environment available in Appendix LABEL:appendix:full-plots.
@@
As baselines, we consider PPO (vectorized cleanRL implementation; Huang et al., 2022) and DeepMDPs (Gelada et al., 2019). Essentially, DeepMDPs are principled auxiliary tasks (the losses $L_{{R}},L_{P}$ presented in Sect. 5) that can be plugged to any RL algorithm to improve the representation learned (with guarantees). The main difference with DeepSPI is that $L_{{R}}$, $L_{P}$ are able to push the updated policy out of the neighborhood by learning the representation via the additional losses, for which updates are not constrained. This means that none of the guarantees presented in this paper apply to DeepMDPs. For a fair comparison, we plugged the DeepMDP losses to (vectorized) PPO and we use the same latent space, network architectures, and distributions as DeepSPI. We use the default cleanRL’s hyperparameters for the three algorithms, except for the data collection ($128$ environments with a horizon of $8$ steps).
@@
As latent space, we use the raw 33D representation obtained after the convolution layers (as recommended by Gelada et al., 2019). For modeling the transition function, we found best to use a mixture of multivariate normal distributions (the transition network outputs $5$ means/diagonal matrices). To deal with the Lipschitz constraints that need to be enforced on the reward and transition functions, we found the most efficient to model $\overline{\!{{R}}},\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ via Lipschitz networks (precisely, we use norm-constrained GroupSort architectures to enforce $1$-Lipschitzness; Anil et al., 2019).
As reported in Fig. 3, DeepSPI provides solid performance and improves on PPO across the majority of environments while being additionally equipped with representation guarantees. Fig. 4 depicts the learning performance of the different algorithms; DeepSPI indeed performs best across the full ALE 57 benchmark suite.
@@
Beyond pure performance, we want to assess whether the world model, learned via DeepSPI, exhibits accurate dynamics. Fig. 5 reports $L_{P}$, $L_{{R}}$ during training. Note that DeepSPI consistently achieves lower transition losses, indicating more accurate transition functions. In contrast to the off-policy setting where Gelada et al. (2019) reported competing transition and reward losses, we did not observe such behavior in our parallel on-policy setting. We attribute this stability to the fact that our losses are always computed under the current policy, unlike off-policy methods that rely on replay buffers.
To probe the predictive quality of the latent model and illustrate Thm. 3, we designed DreamSPI, a naïve variant where DeepSPI learns the world model and representation, while PPO updates the policy from imagined trajectories (Appendix LABEL:appendix:dream-spi). Unlike off-policy methods that update world model and policy from replay buffers, our setting is fully on-policy, making world-model learning and planning more difficult. As a result, the median ALE score of DreamSPI is below DeepSPI and the baselines, though it still learns in several environments and exhibits meaningful behaviours (cf. Fig. 6 & Appendix LABEL:appendix:full-plots). This outcome is not surprising: planning in an on-policy learned model is inherently difficult, and matching direct environment interaction remains challenging. Nevertheless, the value of maintaining a latent model goes well beyond raw scores, as it enables applications in safety, verification, reactive synthesis, and explainability, which we leave for future work.
@@ Appendix A Remark on value functions and episodic processes
$V^{\pi}\mathopen{}\mathclose{{\left(s}}\right)$ $=\begin{cases}\displaystyle\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\mathopen{}\mathclose{{\left(\prod_{i=1}^{t}\mathds{1}\left\{s_{i}\neq s_{\textit{reset}}\right\}\cdot\gamma}}\right){R}\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)\,\Big|\,s_{0}=s\right]&\text{if }s\neq s_{\textit{reset}}\\ 0&\text{otherwise}.\end{cases}$
@@ Appendix B Policy Improvements through Mirror Learning and Convergence Guarantees
Henceforth, fix a policy $\pi\in\Pi$. When taking the supremum, infimum, maximum, or minimum value over states and actions, we always consider actions to be taken from the support of the baseline policy (in the denominator of the quotient).
@@
$\mathcal{N}^{C}$ is lower hemicontinuous (lhc) if, for every policy $\pi$, sequence $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ with $\pi_{n}\to\pi$, and policy $\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)$, there exists a sequence $\mathopen{}\mathclose{{\left(\pi^{\prime}_{n}}}\right)_{n\geq 0}$ with $\pi^{\prime}_{n}\to\pi^{\prime}$ and such that there is a $n_{0}\geq 0$ from which, for all $n\geq n_{0}$, $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$. Therefore, let $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ be a sequence of policies so that $\pi_{n}\to\pi$ and $\pi^{\prime}\in\mathcal{N}\mathopen{}\mathclose{{\left(\pi}}\right)$. Since $\pi_{n}\to\pi$, we have
@@ Appendix D Remark on Safe Policy Improvement Methods
Standard principled safe policy improvement methods (SPI; Thomas et al., 2015; Ghavamzadeh et al., 2016a; Laroche et al., 2019; Simão et al., 2020; Castellini et al., 2023; Wienhöft et al., 2023) do not consider representation learning. Instead, SPI methods assume $\overline{\!{\mathcal{\mathcal{S}}}}\coloneqq\mathcal{S}$ and learn $\overline{\!{{R}}}$, $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ by maximum likelihood estimation with respect to the experience stored in $\mathcal{B}$ collected by the baseline $\pi_{\text{b}}$. Then, the policy improvement relies on finding the best policy in $\overline{\!{\mathcal{M}}}$ that is (probably approximately correctly) guaranteed to improves on the baseline policy (up to an error term $\zeta>0$) against a set of all admissible MDPs, called robust MDPs (Iyengar, 2005; Nilim and Ghaoui, 2005; Wiesemann et al., 2013; Ghavamzadeh et al., 2016b; Suilen et al., 2024):

### v2

@@ 7 Experiments
Figure 5: Aggregate results on stochastic versions of the standard 57 environments from ALE, with 95% confidence intervals (CIs). Higher values for the mean, median, and interquartile mean (IQM) indicate better performance, while a lower optimality gap is preferable (cf. Agarwal et al. 2021b). CIs are obtained through percentile bootstrapping with stratified resampling. Plots per environment available in Appendix LABEL:appendix:full-plots.
Figure 6: Sample efficiency w.r.t. IQM normalized scores on the stochastic ALE-57. Shaded regions give pointwise 95% CIs obtained via percentile stratified bootstrap.
@@
In this section, we evaluate the practical performance of DeepSPI in environments where (i) representation learning is essential and (ii) dynamics are complex. We use the Atari Arcade Learning Environment (ALE; Bellemare et al. 2013) and consider each state as four stacked frames. ALE domains feature a wide range of dynamics; to further introduce stochasticity, we follow Machado et al. (2018) and employ two standard tricks: sticky actions, where with probability $p_{a}$ the previous action is repeated (simulating joystick or reaction-time imperfections), and random initialisation, where the agent begins after $n_{\text{NOOP}}$ initial no-op frames. We set $p_{a}=0.3$ and $n_{\text{NOOP}}=60$.
@@
As baselines, we consider PPO (vectorized cleanRL implementation; Huang et al., 2022) and DeepMDPs (Gelada et al., 2019). Essentially, DeepMDPs are principled auxiliary tasks (the losses $L_{{R}},L_{P}$ presented in Sect. 5) that can be plugged to any RL algorithm to improve the representation learned (with guarantees). The main difference with DeepSPI is that $L_{{R}}$, $L_{P}$ are able to push the updated policy out of the neighborhood by learning the representation via the additional losses, for which updates are not constrained. This means that none of the SPI guarantees presented in this paper apply to DeepMDPs. For a fair comparison, we plugged the DeepMDP losses to (vectorized) PPO, and we use the architecture as for DeepSPI. We use the default cleanRL’s hyperparameters for the three algorithms, except for the data collection ($128$ environments with a horizon of $8$ steps).
@@
As latent space, we use the raw 33D representation obtained after the convolution layers (as recommended and used by Gelada et al., 2019). For modeling the transition function, we found best to use a mixture of multivariate normal distributions (the transition network outputs $5$ means/diagonal matrices). To deal with the Lipschitz constraints that need to be enforced on the reward and transition functions, we found the most efficient to model $\overline{\!{{R}}},\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ via Lipschitz networks (precisely, we use norm-constrained GroupSort architectures to enforce $1$-Lipschitzness; Anil et al., 2019).
As shown in Fig. 5 and 6, DeepSPI delivers strong performance, improving on both PPO and DeepMDP. Notably, these results are obtained while preserving SPI-style properties; a valuable combination, as such theoretical control typically comes at the expense of performance and substantial data requirements. Beyond pure performance, we want to assess whether the world model, learned via DeepSPI, exhibits accurate dynamics.
@@
Fig. 7 reports $L_{P}$, $L_{{R}}$ during training. Note that DeepSPI consistently achieves lower transition losses, indicating more accurate transition functions. We discuss the statistical significance of that statement in Appendix LABEL:appendix:full-plots. In contrast to the off-policy setting where Gelada et al. (2019) reported competing transition and reward losses, we did not observe such behavior in our parallel on-policy setting. We attribute this stability to the fact that our losses are always computed under the current policy, unlike off-policy methods that rely on replay buffers.
To probe the predictive quality of the latent model and illustrate Thm. 3, we introduced DreamSPI, a naïve variant where DeepSPI learns the world model and representation, and PPO updates the policy from imagined trajectories (Appendix LABEL:appendix:dream-spi). Unlike off-policy approaches that exploit replay buffers and update the model at every interaction step, our fully on-policy setting updates the world model only from fresh interaction data, which makes combining model learning and planning inherently more challenging. Even so, DreamSPI achieves learning progress in several environments and exhibits coherent behaviours (cf. Fig. 8 & Appendix LABEL:appendix:full-plots). While its aggregate median score remains below the baselines, this is somehow expected given the stricter data requirements compared to usual model-based approaches. Importantly, the ability to maintain a model offers benefits that extend far beyond raw scores, enabling future applications in safety, verification, and reactive synthesis.
@@ Appendix A Remark on value functions and episodic processes
$V^{\pi}\mathopen{}\mathclose{{\left(s}}\right)$ $=\begin{cases}\displaystyle\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\mathopen{}\mathclose{{\left(\prod_{i=1}^{t}\mathds{1}\left\{s_{i}\neq s_{\textit{reset}}\right\}\cdot\gamma}}\right){R}\mathopen{}\mathclose{{\left(s_{t},a_{t}}}\right)\,\Big|\,s_{0}=s\right]&\text{if }s\neq s_{\textit{reset}}\\ 0&\text{otherwise};\end{cases}$
@@ Appendix B Policy Improvements through Mirror Learning and Convergence Guarantees
Henceforth, fix a policy $\pi\in\Pi$. When taking the supremum, infimum, maximum, or minimum value over states and actions, we always consider actions to be taken from the support of the behavioral policy (in the denominator of the quotient).
@@
$\mathcal{N}^{C}$ is lower hemicontinuous (lhc) if, for every policy $\pi$, sequence $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ with $\pi_{n}\to\pi$, and policy $\pi^{\prime}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi}}\right)$, there exists a sequence $\mathopen{}\mathclose{{\left(\pi^{\prime}_{n}}}\right)_{n\geq 0}$ with $\pi^{\prime}_{n}\to\pi^{\prime}$ and such that there is a $n_{0}\geq 0$ from which, for all $n\geq n_{0}$, $\pi^{\prime}_{n}\in\mathcal{N}^{C}\mathopen{}\mathclose{{\left(\pi_{n}}}\right)$.
Therefore, let $\mathopen{}\mathclose{{\left(\pi_{n}}}\right)_{n\geq 0}$ be a sequence of policies so that $\pi_{n}\to\pi$ and $\pi^{\prime}\in\mathcal{N}\mathopen{}\mathclose{{\left(\pi}}\right)$. Since $\pi_{n}\to\pi$, we have
@@ Appendix D Remark on Safe Policy Improvement Methods
Standard principled safe policy improvement methods (SPI; Thomas et al., 2015; Ghavamzadeh et al., 2016a; Laroche et al., 2019; Simão et al., 2020; Castellini et al., 2023; Wienhöft et al., 2023) do not consider representation learning. Instead, SPI methods assume $\overline{\!{\mathcal{\mathcal{S}}}}\coloneqq\mathcal{S}$ and learn $\overline{\!{{R}}}$, $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ by maximum likelihood estimation with respect to the experience stored in $\mathcal{B}$ collected by the behavioral $\pi_{\text{b}}$. Then, the policy improvement relies on finding the best policy in $\overline{\!{\mathcal{M}}}$ that is (probably approximately correctly) guaranteed to improves on the behavioral policy (up to an error term $\zeta>0$) against a set of all admissible MDPs, called robust MDPs (Iyengar, 2005; Nilim and Ghaoui, 2005; Wiesemann et al., 2013; Ghavamzadeh et al., 2016b; Suilen et al., 2024):